OpenStudy (anonymous):

Find all the points having an x-coordinate of 3 whose distance from the point (-2, -1) is 13.

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

sorry trying to figure this out :)

OpenStudy (skullpatrol):

Start with a sketch please.

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):



OpenStudy (anonymous):

I was told to use the distance formula and the 2 points are (3,y)

OpenStudy (anonymous):

wouldn't it be x -2 and y -1

OpenStudy (anonymous):

ok :) d = √(x2-x1)^2+(y2-y1)^2 would I use my distance formula?

OpenStudy (skullpatrol):

Yes :) now just substitute the values in from your diagram.

OpenStudy (anonymous):

so it would be like this 13=√(x2-(-2))^2+(y2-(-1))^2 :)

OpenStudy (skullpatrol):

looks good

OpenStudy (anonymous):

13=x2+y2+3 then x is 3

OpenStudy (skullpatrol):

x2 = 3

OpenStudy (anonymous):

wow this is extremely difficult. wouldn't i be finding y like this y =mx+b

OpenStudy (anonymous):

i already have X which is 3 right? so i would need the y point which is y = mx +b

OpenStudy (saifoo.khan):

13=(3)^2+y2+3 Solve for y.

OpenStudy (anonymous):

this is the distance formula as i had previous right?

OpenStudy (phi):

ok up to here \[ 13=\sqrt{(x2-(-2))^2+(y2-(-1))^2} \] replace x2 with 3, and just rename y2 to y (less typing) \[ 13 = \sqrt{3 +2)^2 + (y+1)^2}\] square both sides 169= 25 + (y+1)^2 can you finish ?

OpenStudy (anonymous):

got it :)

OpenStudy (phi):

you should get an ugly quadratic: y^2 + 2y -143= 0 which you have to factor to find y

OpenStudy (phi):

another way to do this (assuming you are "allowed to" is